# The radius of statistical efficiency

### v1

@@ 5.3 Matrix completion
The problem of matrix completion corresponds to (24), where the ground truth matrix $M_{\star}$ has low rank and the data matrices $X\sim\mathsf{D}_{x}$ are drawn from some discrete distribution on matrices of the form $X=e_{i}e_{j}^{\top}$. We will focus on the simplified setting where $M_{\star}$ is rank one and positive semidefinite. The standard way to write this problem as stochastic optimization is
$\min_{M\in\mathcal{M}}f(M):=\tfrac{1}{2}\mathop{\operatorname{\mathbb{E}}}_{\mathsf{D}}\Big{[}(\langle X,M\rangle-y\big{)}^{2}\Big{]},$ (34)
where =$\mathcal{M}=\{M\succeq 0:\operatorname{rank}M=1\}$ is the set of rank one PSD matrices. In this section, we compute the RSE of the problem in terms of the graph induced by the support of the distribution $\mathsf{D}_{x}$.
@@
$\nabla f(M_{\star})[\Delta,\Delta]=\operatorname{\mathbb{E}}\langle X,\Delta\rangle^{2}\qquad\forall\Delta\in\bm{S}^{d}.$ (35)
In particular, the optimal Lagrange multipliers for $M_{\star}$ are zero. Forming the factorization $M_{\star}={\beta}_{\star}{\beta}_{\star}^{\top}$, the tangent space to $\mathcal{M}$ at $M_{\star}$ can be written as
$\mathcal{T}=\{{\beta}_{\star}v^{\top}+v{\beta}_{\star}^{\top}:v\in\mathbf{R}^{d}\}.$
Moreover, an elementary computation shows that $\|\Delta\|_{F}^{2}/\|{\beta}_{\star}\|^{2}\|v\|^{2}\in[2,4]$. In particular, $\Delta$ is zero if and only if $v$ is zero. Consequently, the set of ill-conditioned distributions takes the form:
$\mathcal{E}=\bigcup_{v\in\mathbf{R}^{d}\setminus\{0\}}\{\mu\in\mathcal{P}_{2}(\bm{S}^{d}):\operatorname{supp}(X)\subset(v{\beta}_{\star}^{\top})^{\perp}\}.$ (36)
Indeed, estimating ${\rm REG}(\mu)$ at any $\mu\in\mathcal{P}_{2}(\bm{S}^{d})\setminus\mathcal{E}$ is straightforward, and is the content of the following lemma; see Appendix G.1 for a proof.
@@
Consider any measure $\mu\in\mathcal{P}_{2}(\bm{S}^{d})\setminus\mathcal{E}$ and define the matrices
$\Phi_{{\beta}_{\star}}=(I\otimes{\beta}_{\star})+({\beta}_{\star}\otimes I)\qquad\textrm{and}\qquad\Sigma_{\mu}=\operatorname{\mathbb{E}}_{\mu}{\operatorname{vec}}(X){\operatorname{vec}}(X)^{\top}.$
Then the estimate holds:
$2\|\beta_{\star}\|^{2}\leq{\rm REG}(\mu)\cdot\lambda_{\min}(\Phi_{\beta_{\star}}^{\top}\Sigma_{\mu}\Phi_{\beta_{\star}})\leq 4\|\beta_{\star}\|^{2}.$ (37)
@@
${P}=(p_{ij})\quad\text{where}\quad p_{ij}={\operatorname{\mathbb{P}}}\big{\{}(\ell,k)=(i,j)\big{\}}={\operatorname{\mathbb{P}}}\big{\{}X_{ij}=1\big{\}}={\operatorname{\mathbb{E}}}\big{[}X_{ij}\big{]}.$
@@
For each $Q\in\mathcal{Q}$, we let $\mu_{Q}\in\mathcal{P}_{2}({\mathbf{R}}^{d\times d})$ be the distribution over canonical matrices $e_{i}e_{j}^{\top}$ where $(i,j)\sim Q$ [^2] With a slight abuse of notation we use $Q$ to denote both the matrix and distribution over indices. Thus, the set of ill-conditioned distributions encoding matrix completion problems is
$\mathcal{E}^{\text{mc}}:=\big{\{}\mu_{Q}\mid Q\in\mathcal{Q}\big{\}}\cap\mathcal{E},$
where $\mathcal{E}$ is defined in (36). We are now ready to study the Wasserstein distance between $\mu$ and $\mathcal{E}^{\textrm{mc}}$. We will see that this distance relies on the combinatorial structure of the observations.
@@
$G^{*}=(V^{*},E^{*})\text{ be the induced graph given by }V^{*}=\operatorname{supp}({\beta}_{\star}),$ (38)
meaning that $E^{*}$ consists of all the edges in $E$ between elements of $V^{*}$. Further, define
$V^{0}=\{i\notin V^{*}\mid\text{for all }j\in V^{*},(i,j)\notin E\}.$ (39)
Thus, $V^{0}$ corresponds to nodes $i$ with $\beta_{i}=0$ that are not connected to any $j$ for which $\beta_{j}\neq 0.$
We will see shortly that the following assumption characterizes the set of well-posed instances of matrix completion.
Assumption 1.
Consider a graph $G=([d],E)$ with $G^{*}$ and $V^{0}$ defined in (38) and (39), respectively. Suppose the following two hold.
@@
    (Non-bipartite) $G^{*}$ has no connected components that are bipartite.
@@
    (No isolated zeros) The set $V^{0}$ is empty.
@@
[…]
We are now ready to show the main result of this section; see Appendix G.2 for a proof.
Theorem 5.8: (RSE for matrix completion) Let ${P}=(p_{ij})$ and $M_{\star}={\beta}_{\star}{\beta}_{\star}^{\top}$ be the data of the matrix completion problem (34), and let $\mu$ be the distribution of $X$ induced by ${P}.$ Then, $\mu$ is well-posed, i.e., $\mu\notin\mathcal{E}^{{\rm mc}}$, if, and only if, the graph $G$ induced by the problem satisfies Assumption 1. Additionally, the identity holds: ${\rm RSE}(\mu)=\min_{\begin{subarray}{c}A\in\Omega_{{\beta}_{\star}}\\ A\subset\operatorname{supp}({P})\end{subarray}}\,\sum_{ij\in\operatorname{supp}({P})\setminus A}p_{ij}.$ (40) Moreover, computing ${\rm RSE}(\mu)$ is NP-hard in general.
Thus the theorem shows that one can compute ${\rm RSE}(\mu)$ by enumerating over all exceptional edge sets $A\subset\operatorname{supp}({P})$, meaning that $G^{*}_{A}$ either has a connected bipartite component or $V^{0}$ is nonempty. Then, the set $A$ for which the mass $\sum_{ij\in\operatorname{supp}({P})\setminus A}p_{ij}$ is smallest yields ${\rm RSE}(\mu)$. Interestingly, computing ${\rm RSE}(\mu)$ is NP-hard in general, as we show by a reduction from MAXCUT.
@@ 6 Conclusion
In this work, we introduced a new measure of robustness—radius of statistical efficiency (RSE)—for problems of statistical inference and estimation. We computed RSE for a number of test-bed problems, including principal component analysis, generalized linear models, phase retrieval, bilinear sensing, and matrix completion. In all cases, we verified a precise reciprocal relationship between RSE and the intrinsic complexity/sensitivity of the problem instance, thereby paralleling the classical Eckart–Young theorem and its numerous extensions in numerical analysis and optimization. More generally, we obtained sufficient conditions for such a relationship to hold that depend only on local information (gradients, Hessians), rather than an explicit description of the set of ill-conditioned distributions. We believe that this work provides an intriguing new perspective on the interplay between problem difficulty, solution sensitivity, and robustness in statistical inference and learning.
@@ Acknowledgments
We thank John Duchi, Jorge Garza-Vargas, Zaid Harchaoui, and Eitan Levin for insightful conversations during the development of this work.
@@ Appendix A Geometry of the Wasserstein space and distance estimation
In this section, we introduce the necessary background of the Wasserstein geometry and prove a number of results that may be on independent interest. In the following section, we will use many of these results to prove estimates on RSE announced in the paper. We follow standard notation of optimal transport, as set out for example in the monographs of Villani [66] and Santambrogio [59].
Let $(\mathcal{X},\mathsf{d})$ be a separable complete metric space equipped with its Borel $\sigma$-algebra $\mathcal{B}_{\mathcal{X}}$. The primary example for us will be $\mathbf{R}^{d}$ equipped with the $\ell_{p}$ norm. The distance of a point $x\in\mathcal{X}$ to a set $\mathcal{K}\subset\mathcal{X}$ will be denoted by $\operatorname{dist}(x,\mathcal{K})=\inf_{x^{\prime}\in\mathcal{K}}\mathsf{d}(x,x^{\prime})$. The set of Borel probability measures on $\mathcal{X}$ is denoted by $\mathcal{P}(\mathcal{X})$, and will be abbreviated as $\mathcal{P}$ if the space $\mathcal{X}$ is clear from context. The support of a measure $\mu\in\mathcal{P}$, written as $\operatorname{supp}(\mu)$, is the smallest closed set $C\subset\mathcal{X}$ such that the complement $\mathcal{X}\setminus C$ is of zero $\mu$-measure. For any measurable map $T\colon(\Omega,\mathcal{F},\operatorname{\mathbb{P}})\rightarrow(\mathcal{X},\mathcal{B}_{\mathcal{X}})$, the pushforward measure $T_{\#}\mu$ is defined to be
$(T_{\#}\mu)(B)=\mu(T^{-1}(B))\qquad\text{for all }B\in\mathcal{B}_{\mathcal{X}}.$
The support of a random variable on $\mathcal{X}$ is the support of its distribution.
For any $p\geq 1$, the symbol $\mathcal{P}_{p}$ denotes the set of all distributions $\mu$ on $\mathcal{X}$ with finite $p^{\text{th}}$ moment, meaning $\operatorname{\mathbb{E}}_{x\sim\mu}\mathsf{d}(x,x_{0})^{p}<\infty$ for some (and hence any) $x_{0}\in\mathcal{X}$. The Wasserstein-$p$ distance between two measures $\mu,\nu\in\mathcal{P}_{p}$ is defined by:
@@
Here, the set $\Pi(\mu,\nu)$ consists of couplings between $\mu$ and $\nu$, i.e., distributions in $\mathcal{P}_{p}(\mathcal{X}\times\mathcal{X})$ having $\mu$ and $\nu$ as its first and second marginals. An important fact is that the pair $(\mathcal{P}_{p},W_{p})$ is a separable complete metric space in its own right and is called the Wasserstein-$p$ space on $\mathcal{X}$.
We will need a few basic estimates on the $W_{p}$ distance. First, consider any measures $\mu,\nu\in\mathcal{P}_{p}(\mathcal{X})$ and a measurable map $T\colon\mathcal{X}\to\mathcal{X}$ satisfying $\nu=T_{\#}\mu$. Then the law of the random variable $(x,T(x))$ is a coupling between $\mu$ and $\nu$ and therefore the estimate holds:
@@
Another useful observation is that for any measures $\mu,\nu\in\mathcal{P}_{p}$ such that the support of $\nu$ is contained in a set $C\subset\mathcal{K}$, the estimate holds:
$W_{p}^{p}(\mu,\nu)\geq\mathop{\operatorname{\mathbb{E}}}_{x\sim\mu}\operatorname{dist}^{p}(x,C).$ (42)
Consequently, equality holds in (41) when $T$ is a projection, the content of the following lemma.
@@
Consider a measure $\mu\in\mathcal{P}_{p}$ and a set $\mathcal{K}\subset\mathcal{X}$. Suppose that the metric projection $\mathsf{P}_{\mathcal{K}}$ admits a measurable selection $s\colon\mathcal{X}\rightarrow\mathcal{X}$. Then equality holds:
@@
and therefore $\nu$ lies in $\mathcal{P}_{p}$. Next, taking into account that the support of $\nu$ is contained in $\mathcal{K}$, combining (41) and (42) completes the proof. ∎
[…]
@@
Given a set of measures $\mathcal{V}\subset\mathcal{P}_{p}$, the distance function $W_{2}(\cdot,\mathcal{V})$ is defined in the usual way
$W_{2}(\mu,\mathcal{V})=\inf_{\nu\in\mathcal{V}}~{}W_{2}(\mu,\nu).$

### v2

@@ 5.3 Matrix completion
The problem of matrix completion corresponds to the measurement model (26) where the ground truth matrix $M_{\star}$ has low rank and the feature data matrices $X\sim\mathsf{D}_{X}$ are drawn from a certain discrete distribution. We will focus on the simplified setting where $M_{\star}$ is rank-one and positive semidefinite. The standard way to write this problem as stochastic optimization is
$\min_{M\in\mathcal{M}}f(M):=\mathop{\operatorname{\mathbb{E}}}_{(X,y)\sim\mathsf{D}}\tfrac{1}{2}\big(\langle X,M\rangle-y\big)^{2},$ (36)
where $\mathcal{M}:=\{M\in\bm{S}_{+}^{d}\mid\operatorname{rank}(M)=1\}$ is the manifold of PSD rank-one $d\times d$ matrices. In this section, we will compute the RSE for matrix completion problems of the form (36) in terms of the graph induced by the support of $\mathsf{D}_{X}$.
@@
$\nabla^{2}f(M_{\star})[\Delta,\Delta]=\operatorname{\mathbb{E}}\langle X,\Delta\rangle^{2}\qquad\forall\Delta\in\mathbf{R}^{d\times d}.$ (37)
In particular, the optimal Lagrange multipliers at $M_{\star}$ are zero. Forming the factorization $M_{\star}={\beta}_{\star}{\beta}_{\star}^{\top}$, the tangent space of $\mathcal{M}$ at $M_{\star}$ may be written as
$\mathcal{T}=\{{\beta}_{\star}v^{\top}+v{\beta}_{\star}^{\top}\mid v\in\mathbf{R}^{d}\}.$
Moreover, an elementary computation shows $\|\Delta\|_{F}^{2}/\|{\beta}_{\star}\|^{2}_{2}\|v\|^{2}_{2}\in[2,4]$ for all $\Delta={\beta}_{\star}v^{\top}+v{\beta}_{\star}^{\top}$ and nonzero $v\in\mathbf{R}^{d}$. In particular, $\Delta$ is zero if and only if $v$ is zero. Consequently, the set of ill-conditioned distributions for (36) takes the form:
$\mathcal{E}=\bigcup_{v\in\mathbf{R}^{d}\setminus\{0\}}\big\{\mu\in\mathcal{P}_{2}(\mathbf{R}^{d\times d})\mid\operatorname{supp}(\mu)\subset(\beta_{\star}v^{\top}+v{\beta}_{\star}^{\top})^{\perp}\big\}.$ (38)
Estimating ${\rm REG}(\mu)$ at any $\mu\in\mathcal{P}_{2}(\mathbf{R}^{d\times d})\setminus\mathcal{E}$ is straightforward, and is the content of the following lemma; see Appendix G.1 for a proof.
@@
Consider any measure $\mu\in\mathcal{P}_{2}(\mathbf{R}^{d\times d})\setminus\mathcal{E}$ and define the matrices
$\Phi_{{\beta}_{\star}}:=(I\otimes{\beta}_{\star})+({\beta}_{\star}\otimes I)\qquad\textrm{and}\qquad\Sigma_{\mu}:=\operatorname{\mathbb{E}}_{\mu}{\operatorname{vec}}(X){\operatorname{vec}}(X)^{\top}.$
Then for the problem (36), the following estimate holds:
$2\|\beta_{\star}\|^{2}_{2}\leq{\rm REG}(\mu)\cdot\lambda_{\min}(\Phi_{\beta_{\star}}^{\top}\Sigma_{\mu}\Phi_{\beta_{\star}})\leq 4\|\beta_{\star}\|^{2}_{2}.$ (39)
@@
${P}=(p_{ij})\quad\text{where}\quad p_{ij}={\operatorname{\mathbb{P}}}\big\{X=e_{i}e_{j}^{\top}\big\}.$
We denote the set of all such matrices by
[…]
Note that we allow for $\sum_{i,j}p_{ij}<1$, meaning that the zero matrix may be sampled with positive probability, in which case no entries are observed. For each $P=(p_{ij})$ in $\mathcal{Q}$ , we let $\mu_{P}\in\mathcal{P}_{2}(\mathbf{R}^{d\times d})$ denote the corresponding discrete distribution given by
$\mu_{P}(0)=1-\sum_{i,j}p_{ij}\quad\text{ and }\quad\mu_{P}\big(e_{i}e_{j}^{\top}\big)=p_{ij}.$
Then the set of ill-conditioned distributions for matrix completion problems of the form (36) is
$\mathcal{E}^{\text{mc}}:=\{\mu_{P}\mid P\in\mathcal{Q}\}\cap\mathcal{E},$
where $\mathcal{E}$ is given by (38). We are now ready to study the Wasserstein-$2$ distance from $\mu_{P}$ to $\mathcal{E}^{\textrm{mc}}$.
@@
meaning that $E^{*}_{A}$ consists of all the edges in $E_{A}$ between vertices in $V^{*}$, and we define
$V^{0}_{A}:=\big\{i\in\hbox{\set@color\hskip 2.22221pt\hskip-2.22221pt\hbox{\set@color$\lfloor$}\hskip-2.22221pt\hskip-2.22221pt\raisebox{0.0pt}{\hbox{\set@color$\lceil$}}\hskip-2.22221pt\hskip 2.22221pt}d\hbox{\set@color\hskip 2.22221pt\hskip-2.22221pt\hbox{\set@color$\rfloor$}\hskip-2.22221pt\hskip-2.22221pt\raisebox{0.0pt}{\hbox{\set@color$\rceil$}}\hskip-2.22221pt\hskip 2.22221pt}\setminus V^{*}\mid\{i,j\}\notin E_{A}\text{ for all }j\in V^{*}\big\}.$ (41)
Thus, $V^{0}_{A}$ is comprised of the vertices $i$ with ${\beta_{\star}}_{i}=0$ that are not adjacent in $G_{A}$ to any vertex $j$ for which ${\beta_{\star}}_{j}\neq 0$.
With the preceding notation in hand, we may now describe precisely the combinatorial structure that characterizes well-posed instances of matrix completion.
Assumption 1.
@@
    (No bipartite components of $G_{A}^{*}$) Every connected component of $G^{*}_{A}$ is non-bipartite.
@@
    (No isolated zeros of $\beta$$_{\star}$) The set $V^{0}_{A}$ is empty.
We will see that the set
@@
plays a crucial role in computing ${\rm RSE}(\mu_{P})$. We are now ready to state the main result of this section; see Appendix G.2 for a proof.
Theorem 5.8: (RSE for matrix completion) Let ${P}=(p_{ij})$ and $M_{\star}={\beta}_{\star}{\beta}_{\star}^{\top}$ be the data for a matrix completion problem of the form (36) with $X\sim\mu_{P}$. Then $\mu_{P}$ is well-posed, i.e., $\mu_{P}\notin\mathcal{E}^{{\rm mc}}$, if and only if the graph $G_{\operatorname{supp}(P)}$ satisfies Assumption 1, i.e., $\operatorname{supp}(P)\notin\Omega_{{\beta}_{\star}}$. Additionally, the following identity holds: ${\rm RSE}(\mu_{P})^{2}=\min_{\begin{subarray}{c}A\in\Omega_{{\beta}_{\star}}\\ A\subset\operatorname{supp}(P)\end{subarray}}\sum_{(i,j)\in\operatorname{supp}({P})\setminus A}p_{ij}.$ (42) Moreover, computing ${\rm RSE}(\mu_{P})$ is NP-hard in general.
Thus, Theorem 5.3 shows
$\mathcal{E}^{\text{mc}}=\big\{\mu_{P}\mid P\in\mathcal{Q}~\text{and}~{\operatorname{supp}}(P)\in\Omega_{{\beta}_{\star}}\big\}$ (43)
and that one can compute ${\rm RSE}(\mu_{P})$ by enumerating over all symmetric subsets $A\subset\operatorname{supp}({P})$ such that either $G^{*}_{A}$ has a bipartite connected component or $V^{0}_{A}$ is nonempty. Then any such subset $A$ for which the mass $\sum_{(i,j)\in\operatorname{supp}({P})\setminus A}p_{ij}$ is smallest yields ${\rm RSE}(\mu_{P})^{2}$. Interestingly, computing ${\rm RSE}(\mu_{P})$ is NP-hard in general, which follows from a polynomial-time reduction from the MaxCut problem.
@@ 6 Conclusion
In this work, we introduced a new measure of robustness—radius of statistical efficiency (RSE)—for problems of statistical inference and estimation. We computed RSE for a number of testbed problems, including principal component analysis, generalized linear models, phase retrieval, bilinear sensing, and matrix completion. In all cases, we verified a precise reciprocal relationship between RSE and the intrinsic complexity/sensitivity of the problem instance, thereby paralleling the classical Eckart–Young theorem and its numerous extensions in numerical analysis and optimization. More generally, we obtained sufficient conditions for such a relationship to hold that depend only on local information (gradients, Hessians), rather than an explicit description of the set of ill-conditioned distributions. We believe that this work provides an intriguing new perspective on the interplay between problem difficulty, solution sensitivity, and robustness in statistical inference and learning.
@@ Acknowledgments
We thank John Duchi, Jorge Garza-Vargas, Zaid Harchaoui, and Eitan Levin for insightful conversations during the development of this work. We also thank the anonymous reviewers for their useful feedback, especially for their improvement to the proof of Lemma A.13.
@@ Appendix A Geometry of the Wasserstein space and distance estimation
In this section, we introduce the necessary background of the Wasserstein geometry and prove a number of results that may be of independent interest; in the subsequent section, we will use many of these results to prove estimates on RSE announced in the paper. We follow standard notation of optimal transport, as set out for example in the monographs of Villani [70] and Santambrogio [62].
Let $(\mathcal{X},\mathsf{d})$ be a separable complete metric space equipped with its Borel $\sigma$-algebra $\mathcal{B}_{\mathcal{X}}$. The primary example for us will be a Euclidean space $(\mathbf{E},\|\cdot\|)$, specifically $\mathbf{R}^{d}$ or $\mathbf{R}^{m\times n}$. The distance of a point $x\in\mathcal{X}$ to a set $\mathcal{K}\subset\mathcal{X}$ will be denoted by $\operatorname{dist}(x,\mathcal{K})=\inf_{x^{\prime}\in\mathcal{K}}\mathsf{d}(x,x^{\prime})$. The set of Borel probability measures on $\mathcal{X}$ is denoted by $\mathcal{P}(\mathcal{X})$, and will be abbreviated as $\mathcal{P}$ if the space $\mathcal{X}$ is clear from context. The support of a measure $\mu\in\mathcal{P}$, written as $\operatorname{supp}(\mu)$, is the smallest closed set $C\subset\mathcal{X}$ such that the complement $\mathcal{X}\setminus C$ is of zero $\mu$-measure. For any measurable map $T\colon(\Omega,\mathcal{F},\operatorname{\mathbb{P}})\rightarrow(\mathcal{X},\mathcal{B}_{\mathcal{X}})$, the pushforward measure $T_{\#}\operatorname{\mathbb{P}}$ is defined by setting
$(T_{\#}\operatorname{\mathbb{P}})(B)=\operatorname{\mathbb{P}}(T^{-1}(B))\qquad\text{for all }B\in\mathcal{B}_{\mathcal{X}}.$
The support of the random variable $T$ in $\mathcal{X}$ is the support of its distribution $T_{\#}\operatorname{\mathbb{P}}$.
For any $p\geq 1$, the symbol $\mathcal{P}_{p}$ denotes the set of all distributions $\mu$ on $\mathcal{X}$ with finite $p^{\text{th}}$ moment, meaning $\operatorname{\mathbb{E}}_{x\sim\mu}\mathsf{d}(x,x_{0})^{p}<\infty$ for some (and hence any) $x_{0}\in\mathcal{X}$. The Wasserstein-$p$ distance between two measures $\mu,\nu\in\mathcal{P}_{p}$ is given by
@@
Here, the set $\Pi(\mu,\nu)$ is comprised of the couplings between $\mu$ and $\nu$, i.e., distributions on $\mathcal{X}\times\mathcal{X}$ having $\mu$ and $\nu$ as its first and second marginals. An important fact is that the pair $(\mathcal{P}_{p},W_{p})$ is a separable complete metric space in its own right and is called the Wasserstein-$p$ space on $\mathcal{X}$.
We will need a few basic estimates on the $W_{p}$ distance. First, consider any measures $\mu,\nu\in\mathcal{P}_{p}$ and a measurable map $T\colon\mathcal{X}\to\mathcal{X}$ satisfying $\nu=T_{\#}\mu$. Then the law of the random variable $(x,T(x))$, with $x\sim\mu$, is a coupling between $\mu$ and $\nu$ and therefore the estimate holds:
@@
Another useful observation is that for any measures $\mu,\nu\in\mathcal{P}_{p}$ such that the support of $\nu$ is contained in the closure of a set $\mathcal{K}\subset\mathcal{X}$, the estimate holds:
$W_{p}^{p}(\mu,\nu)\geq\mathop{\operatorname{\mathbb{E}}}_{x\sim\mu}\operatorname{dist}(x,\mathcal{K})^{p}.$ (45)
Consequently, equality holds in (44) when $T$ is a projection—the content of the following lemma.
@@
Consider a measure $\mu\in\mathcal{P}_{p}$ and a set $\mathcal{K}\subset\mathcal{X}$. Suppose that the metric projection $\mathsf{P}_{\mathcal{K}}$ admits a measurable selection $s\colon\mathcal{X}\rightarrow\mathcal{K}$. Then equality holds:
@@
and therefore $\nu$ lies in $\mathcal{P}_{p}$. Next, taking into account that the support of $\nu$ is contained in the closure of $\mathcal{K}$, combining (44) and (45) completes the proof. ∎
For the rest of the section, we will focus exclusively on the setting where $\mathcal{X}$ is the standard $d$-dimensional Euclidean space $\mathbf{R}^{d}$ equipped with the dot product $\langle x,y\rangle=x^{\top}y$ and the induced norm $\|x\|_{2}=\sqrt{\langle x,x\rangle}$. We denote the second-moment matrix of any measure $\mu\in\mathcal{P}_{2}$ by the symbol
@@
Given a set of measures $\mathcal{V}\subset\mathcal{P}_{2}$, the distance function $W_{2}(\cdot,\mathcal{V})$ is defined in the usual way:
$W_{2}(\mu,\mathcal{V})=\inf_{\nu\in\mathcal{V}}\,W_{2}(\mu,\nu).$